# Geometric Optimal Control of Mechanical Systems with Gravitational and Resistive Force

### v1

@@ IV Geometric Intuition
In this section, we describe the geometric meaning of the optimal spline equation. Even though the optimal control equation in 13 is general for the external force, we narrow down the scope for our interest in the gravitational and drag forces. We consider a simple model for a resistive force and include a potential force,
$F_{\text{ext}}=-d\mathcal{V}-D({q})\dot{{q}},$ (15)
where $D$ is a drag tensor containing information about the drag coefficient for each direction with respect to a configuration, $\mathcal{V}:Q\to\mathbb{R}$ is the potential energy as a smooth function of a configuration (i.e., a smooth scalar field) on a configuration manifold $Q$, and $d\mathcal{V}$ is an Euclidean gradient of potential energy. With this setting, the optimal control system of interest is defined by the data $(Q,M,N,a_{\text{ext}})\equiv(Q,M,N,\mathcal{V},D)$.
For a straightforward explanation through this section, we consider the acceleration-based norm cost, which makes the dual inertia metric the cometric so that the data we consider is reduced to $(Q,M,\mathcal{V},D)$,
$N=M\text{ and }\lVert u\rVert^{2}_{N}=\lVert u\rVert^{2}_{M}$ (16)
@@ IV-A Inertial effect
Consider Without the external force and with the acceleration-based cost, the optimal spline equation in 13 reduces to a Riemannian spline equation governed by the Riemannian curvature tensor term,
$\nabla_{v}^{2}u+R\left(u,v\right)v=0.$ (17)
The Riemannian spline equation resembles the geodesic variation (Jacobi) equation [15], allowing for a similar geometric interpretation. On flat manifolds with zero curvature, as illustrated in Fig. 2(a), control input $u$ increases or decreases linearly along the trajectory. On manifolds of positive curvature, geodesics tend to converge, and the Riemannian curvature term, as illustrated in Fig. 2(b), causes control inputs to point inward, leading to trajectory convergence. Conversely, on manifolds of negative curvature, control inputs point outward, resulting in trajectory divergence, as illustrated in Fig. 2(c).
IV-B Potential force
Consider a conservative dynamical system, which means that the external force only includes a potential force. The equivalent acceleration generated by the potential force is a Riemannian gradient of potential energy with respect to an inertia metric $M$: $\text{grad}\equiv\text{grad}_{M}$ [15],
$-\text{grad}{}(\mathcal{V})=-\left\langle d\mathcal{V},\bullet\right\rangle_{M^{*}}=a_{\text{ext}}.$ (18)
@@ IV-B Potential force
Figure 5: (a) Optimal trajectories for a UR5 moving its first three joints under the combinational effect of inertia, gravitation, and drag force. Tissot indicatrices visualize the metric space at given points of the trajectory. The red-colored curve denotes an optimal trajectory under the torque-based cometric. The blue-colored curve denotes an optimal trajectory under the Euclidean metric, and it is the same as the interpolation results between initial and final conditions. The set of surfaces denotes a level set of a potential field. (b) An optimal trajectory of UR5 manipulator under the torque-based cometric in task space. (c) An optimal trajectory of UR5 manipulator under the Euclidean metric or a trajectory interpolated between initial and final conditions in task space. (d) A time versus a torque norm cost. A cost can be derived by the time-integral of a torque norm.
A Riemannian Hessian of a scalar field (e.g., a potential field) associated with an inertia metric $M$ describes a second-order effect of a function on the manifold: $\text{Hess}\equiv\text{Hess}_{M}$ [15],
$\text{Hess}(\mathcal{V})(X,Y)=\left\langle\nabla_{X}\text{grad}{}(\mathcal{V}),Y\right\rangle_{M}.$ (19)
@@
The potential force term in 13 reduces to an inner product of the acceleration-based control input $u$ and a Riemmanian gradient of potential energy. By the definition of Riemannian Hessian, this term is equivalent to a Riemannian Hessian of potential energy,
$-2\left\langle\alpha\,;\,\nabla_{\bullet}a_{\text{ext}}\right\rangle=\left\langle\nabla_{\bullet}\text{grad}(\mathcal{V}),u\right\rangle_{M}=\text{Hess}(\mathcal{V})(\bullet,u),$ (20)
The potential energy term in optimal spline equation 13 is equivalent to the vector contracting potential Hessian with a control input $u$. Because of the sign difference from the definition of the potential force, a positive local curvature of potential fields makes the control input points outward. From the viewpoint of mechanics, the controller uses the gravity force to reduce the use of the control input as much as possible. Thus, the optimal trajectory prefers the route where the control input aligns with the direction changing in a potential force and avoids the misaligned direction.
Examples: As illustrated in Fig. 3, we compare two optimal trajectories for a two-link manipulator with and without gravity to evaluate its effect. The inertia tensor is derived by pulling back each link’s individual inertia tensor through its Jacobian. A detailed derivation of a mass matrix and a potential field can be found in [17, 18]. For this example, we use a torque norm as the cost functional, with a torque-based cometric tensor that reduces to the identity matrix due to its physical configurations. Two scenarios of optimal control are considered. The initial and final configuration conditions of each scenario are
@@ IV-C Resistive Force
With a drag effect, a trajectory minimizing a cost is affected by three different terms,
$\lVert u\rVert^{2}_{N}$ $=\lVert\nabla_{v}v\rVert^{2}_{N}+\lVert v\rVert^{2}_{D^{\prime}}+2\left\langle\left\langle\nabla_{v}v,\bullet\right\rangle_{M},\left\langle v,\bullet\right\rangle_{D}\right\rangle_{\tilde{M}},$ (22)
where an induced drag metric tensor $D^{\prime}=D^{T}\tilde{M}D$ encodes the anti-dual of a cometric tensor with respect to a drag tensor. The first term is the covariant acceleration norm, dominated by an inertia manifold. The second term can be interpreted as the drag force norm associated with the inertia metric or the velocity norm on the manifold associated with an induced drag metric tensor. To minimize the integral of the velocity norm along the trajectory, a short pathlength on the drag manifold and constant-speed profile is desirable due to this term [13, 19]. The third term is the coupling effect represented by the inner product of inertia and drag forces with respect to the cometric tensor. Without a drag effect, the second and third terms vanish.
@@ V Results
We applied the methods in 13 to construct an optimal control scheme for a UR5 manipulator. The initial and final configuration conditions are $q_{0}=\left[0.1,0.1,\pi/2\right]^{T}$, $q_{T}=\left[\pi/2,-1,1\right]^{T}$ and fixed the remaining joint values as $0.1$. The initial and final velocity conditions are zero velocities. Note that we only give the command to the first three joints starting from the base joint to decrease computational cost and visualize the trajectory, even though the method can be applied to high-dimensional systems. The dynamical system of UR5 includes the gravity and joint friction forces. The mass and linkage length information can be found in [14]. We assume a viscous friction model for simplicity. To visualize the robots, we use MATLAB’s Robotics System Toolbox. To solve the boundary value problems in this paper, we use the multiple shooting method based on the CasADi’s automatic differentiation [20].
As illustrated in Fig. 5, we compare two optimal trajectories for the UR5 manipulator under the same initial and final conditions, accounting for joint friction and gravitational effects. The Euclidean metric trajectory interpolates between the start and end states, while the torque-based cometric trajectory accounts for inertia, resulting in a path that may appear longer in joint space but is more efficient. Fig. 5(b-c) illustrates the differences between the torque-optimal and Euclidean metric trajectories in task space. The torque-optimal trajectory initially folds the arm to minimize inertia effects and then stretches it to reach the final configuration, where inertia has less impact. As illustrated in Fig. 5(d), the cost of the torque-optimal trajectory increases moderately over time compared to the Euclidean metric trajectory and represents a more efficient route overall so that the control input (or the trajectory) aligns with the shorter axis of the Tissot ellipsoids.
@@ VI CONCLUSIONS
In this paper, we have presented a geometric framework for understanding optimal control in mechanical systems, considering the effects of inertia and external forces (e.g., drag and gravitational forces). By employing Riemannian geometry to describe these systems, we gained insights into how these external forces influence optimal trajectories. We derived the optimal control equations based on the Pontryagin Maximum Principle and demonstrated their application to two-link and UR5 robotic manipulators, highlighting the curvature effects of each manifold and the path-shortening influence of resistive forces. This approach not only enhances our understanding of trajectory optimization but also provides a foundation for applying geometric optimal control methods to a wide range of robotic systems (e.g., locomoting systems [21, 22]) and mechanical applications. Future work could explore the integration of additional forces and constraints (e.g., nonholonomic constraints [23]), further broadening the applicability of this geometric approach to more complex systems.
ACKNOWLEDGMENT
We acknowledge the support in part by the Office of Naval Research under awards N00014-23-1-2171 and the Brazilian agency grants CNPq 309847/2021-4, 402320/2023-9 and FAPERJ JCNE E-26/203.262/2017.

### v2

@@ IV Geometric Intuition
In this section, we describe the geometric meaning of the optimal spline equation. Even though the optimal control equation in 11 is general for the external force, we narrow down the scope for our interest in the gravitational and drag forces. We consider a simple model for a resistive force and include a potential force,
$a_{\text{ext}}^{i}=-M^{ij}\left(\partial_{{q}^{j}}\mathcal{V}({q})+D_{jk}({q})\dot{{q}}^{k}\right),$ (13)
where $D$ is a drag tensor containing information about the drag coefficient for each direction with respect to a configuration, $\mathcal{V}:Q\to\mathbb{R}$ is the potential energy as a smooth function of a configuration (i.e., a smooth scalar field) on a configuration manifold $Q$, and $\partial_{{q}^{j}}\mathcal{V}$ is a partial derivative of potential energy along ${q}^{j}$ direction. With this setting, the optimal control system of interest is defined by the data $(Q,M,N,F_{\text{ext}})\equiv(Q,M,N,\mathcal{V},D)$.
For a straightforward explanation through this section, we consider the acceleration-based norm cost, which makes the dual inertia metric so that the data we consider is reduced to $(Q,M,\mathcal{V},D)$,
$N=M\text{ and }u^{i}=\dfrac{1}{2}M^{ij}\alpha_{j}.$ (14)
@@ IV-A Inertial Effect
Consider the situation without the external force and with the acceleration-based cost, the optimal spline equation in 11 reduces to a Riemannian spline equation governed by the Riemannian curvature tensor term,
$\left(\nabla_{\dot{{q}}}^{2}a\right)^{i}+R^{i}_{jkl}\dot{{q}}^{j}\dot{{q}}^{k}a^{l}=0.$ (15)
The Riemannian spline equation resembles the geodesic variation (Jacobi) equation [15], allowing for a similar geometric interpretation. On flat manifolds with zero curvature, control input $u$ increases or decreases linearly along the trajectory. On manifolds of positive curvature, geodesics tend to converge, and the Riemannian curvature term causes control inputs to point inward, leading to trajectory convergence. Conversely, on manifolds of negative curvature, control inputs point outward, resulting in trajectory divergence.
IV-B Potential Force Effect
Consider a conservative dynamical system, which means that the external force only includes a potential force. The equivalent acceleration generated by the potential force is a Riemannian gradient of potential energy with respect to an inertia metric $M$: $\text{grad}\equiv\text{grad}_{M}$  [15, 8],
$-(\text{grad}{}\mathcal{V})^{i}=M^{ij}\partial_{{q}^{j}}\mathcal{V}=a_{\text{ext}}^{i}.$ (16)
@@ IV-B Potential Force Effect
[…]
A Riemannian Hessian of a scalar field (e.g., a potential field) associated with an inertia metric $M$ describes a second-order derivative effect of a scalar field on the manifold: $\text{Hess}\equiv\text{Hess}_{M}$ [15, 8],
$(\text{Hess}\mathcal{V})_{ij}=\partial_{{q}^{i}}\partial_{{q}^{j}}\mathcal{V}-\Gamma^{k}_{ij}\partial_{{q}^{k}}\mathcal{V}=M_{jk}(\nabla_{\partial_{{q}^{i}}}\text{grad}{}\mathcal{V})^{k}.$ (17)
@@
The potential force term in 11 can be simplified using 16 and 17, yielding a Riemannian Hessian of the potential energy:
$\alpha_{k}\left(\nabla_{\partial_{{q}^{i}}}a_{\text{ext}}\right)^{k}=-\alpha_{k}\left(\nabla_{\partial_{{q}^{i}}}\text{grad}{}\mathcal{V}\right)^{k}\\ =-M^{jk}\alpha_{k}(\text{Hess}\mathcal{V})_{ij}=-2u^{j}(\text{Hess}\mathcal{V})_{ij}.$ (18)
With the assumption 14, the potential energy term corresponds to the contraction of the potential Hessian with the control input. Due to the sign difference in the definition of potential force, a positive local curvature of the potential field directs the control input outward. From a mechanical perspective, the controller leverages gravitational force to minimize control effort. Consequently, the optimal trajectory tends to follow paths where the control input aligns with changes in potential force, while avoiding misaligned directions.
Examples: As illustrated in Fig. 2, we compare two optimal trajectories for a two-link manipulator with and without gravity to evaluate its effect. The inertia tensor is derived by pulling back each link’s individual inertia tensor through its Jacobian. A detailed derivation of a mass matrix and a potential field can be found in [17, 18]. For this example, we use a torque norm as the cost functional, with an actuation-based cometric tensor that reduces to the identity matrix due to its physical configurations. Two scenarios of optimal control are considered. The initial and final configuration conditions of each scenario are
@@ IV-C Resistive Force Effect
With drag effects, the cost functional can be decomposed into three distinct terms by combining 6b and 6c:
$\lVert u\rVert_{M}^{2}=M_{ij}u^{i}u^{j}\\ =M_{ij}(\nabla_{\dot{{q}}}\dot{{q}})^{i}(\nabla_{\dot{{q}}}\dot{{q}})^{j}+D^{\prime}_{ij}\dot{{q}}^{i}\dot{{q}}^{j}-2(F_{D})_{j}(\nabla_{\dot{{q}}}\dot{{q}})^{j}.$ (20)
where $(F_{D})_{j}=D_{ij}\dot{{q}}^{i}$ is a drag force covector, and $D^{\prime}=D^{T}MD$ is an induced drag metric tensor.
  • The first term represents the covariant acceleration norm, primarily influenced by the inertia manifold.
  • The second term corresponds to the drag force norm associated with the inertia metric, or equivalently, the velocity norm on the drag manifold. Minimizing the integral of this norm favors shorter paths on the drag manifold and a constant-speed profile [19, 20].
  • The third term captures the coupling effect, given by the product of covariant acceleration and drag forces.
  • Without drag effects, the second and third terms vanish, reducing the formulation to the classical inertia-dominated case.
@@ V Results
We applied the methods in 11 to construct an optimal control scheme for a UR5 manipulator. The initial and final configuration conditions are $q_{0}=\left[0.1,0.1,\pi/2\right]^{T}$, $q_{T}=\left[\pi/2,-1,1\right]^{T}$ and fixed the remaining joint values as $0.1$. The initial and final velocity conditions are zero velocities. Note that we only give the command to the first three joints starting from the base joint to decrease computational cost and visualize the trajectory, even though the method can be applied to high-dimensional systems. The dynamical system of UR5 includes the gravity and joint friction forces. The mass and linkage length information can be found in [14]. We assume a linear friction model for simplicity. To visualize the robots, we use MATLAB’s Robotics System Toolbox. To solve the boundary value problems in this paper, we use the multiple shooting method based on the CasADi’s automatic differentiation [21].
As illustrated in Fig. 4, we compare two optimal trajectories for the UR5 manipulator under the same initial and final conditions, accounting for joint friction and gravitational effects. The Euclidean metric trajectory interpolates between the start and end states, while the actuation-based cometric trajectory accounts for inertia, resulting in a path that may appear longer in joint space but is more efficient. Fig. 4(b-c) illustrates the differences between the actuation-optimal and Euclidean metric trajectories in task space. The actuation-optimal trajectory initially folds the arm to minimize inertia effects and then stretches it to reach the final configuration, where inertia has less impact. As illustrated in Fig. 4(d), the cost of the actuation-optimal trajectory increases moderately over time compared to the Euclidean metric trajectory and represents a more efficient route overall so that the control input (or the trajectory) aligns with the shorter axis of the Tissot ellipsoids.
@@ VI CONCLUSIONS
In this paper, we have presented a geometric framework for understanding optimal motions in mechanical systems, considering the effects of inertia and external forces (e.g., drag and gravitational forces). By employing Riemannian geometry to describe these systems, we gained insights into how these external forces influence optimal trajectories. We derived the optimal control equations based on the Pontryagin Maximum Principle [11, 12] and demonstrated their application to two-link and UR5 robotic manipulators, highlighting the curvature effects of each manifold and the path-shortening influence of resistive forces. This approach not only enhances our understanding of trajectory optimization but also provides a foundation for applying Riemannian variational calculus to a wide range of robotic systems (e.g., locomoting systems [22, 23]) and mechanical applications. Future work could explore the integration of additional forces and constraints (e.g., nonholonomic constraints [4]), further broadening the applicability of this geometric approach to more complex systems.